OpenStudy (anonymous):

Solve by Substitution 10x + 3y = 8 2x + y = 2 I need the steps please.

OpenStudy (anonymous):

I have the answer x = 1/2 y = 1

OpenStudy (anonymous):

i've given you the steps before

OpenStudy (anonymous):

this -> from the second equation 2x+y = 2 y = 2-2x then substitute y to the first equation 10x+3y=8 10x+3(2-2x) = 8 10x+6-6x=8 10x-6x=8-6 4x=2 x=0.5 tell me which part you dont understand

OpenStudy (anonymous):

the answer I found the x =1/2 and y =1

OpenStudy (anonymous):

yes, that's the same as my answer x=0.5 0.5 is the decimal form of 1/2

OpenStudy (anonymous):

now I get it thanks a lot.

OpenStudy (anonymous):

you're welcome